OpenStudy (anonymous):

Solve x – 5 / x – 1 = x + 3 / x 8 0 -1 -8

Parth (parthkohli):

\[\dfrac{x - 5}{x - 1} = \dfrac{x + 3}{x}\]Multiply both sides by \(x\) and then \(x - 1\). In other words, just cross multiply.\[x(x - 5) = (x - 1)(x + 3)\]Expand both sides and you will get a quadratic equation.

OpenStudy (anonymous):

srry parth im still clueless

Parth (parthkohli):

OK, well. Do you know how to simplify \(x(x - 5)\)?

OpenStudy (anonymous):

no

Parth (parthkohli):

I suggest you to read your book/watch the lecture again... very closely.